OpenStudy (anonymous):

find the area inside the oval limacon r=6+2cos(theta)

OpenStudy (amistre64):

polar integral eh

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

trying to recall:\[\int\int r(t)dtdr\]or is that going to far?

OpenStudy (anonymous):

the formula is \[\int\limits_{a}^{b}1/2 (r(\theta))^2 d \theta\]

OpenStudy (amistre64):

\[\int\limits_{a}^{b}\frac12 (r(\theta))^2 d \theta\] by chance?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

and t is from 0 to 2pi; and they give you r(t) \[2*\int\limits_{0}^{pi}\frac12 (6+2cos(t)))^2 dt\] \[2*\int\limits_{0}^{pi}\frac12 4cos^2(t)+24cos(t)+36 dt\] \[2*\int\limits_{0}^{pi} 2cos^2(t)+12cos(t)+18 dt\] might have to reduce power on the cos^2

OpenStudy (amistre64):

\[cos(2t)=cos^2(t)-sin^2(t)\] \[cos(2t)=cos^2(t)-(1-cos^2(t))\] \[cos(2t)=cos^2(t)-1+cos^2(t))\] \[cos(2t)=2cos^2(t)-1\] \[\frac{1+cos(2t)}{2}=cos^2(t)\]

OpenStudy (amistre64):

im curious what the 1/2 signifies in there

OpenStudy (amistre64):

i think my doubling it was premature

OpenStudy (amistre64):

would you agree?

OpenStudy (amistre64):

but then, im only integrating the top half, so i would need to double it ...

OpenStudy (amistre64):

\[2*\int\limits_{0}^{pi} 2\frac{1+cos(2t)}{2}+12cos(t)+18 dt\] \[2*\int\limits_{0}^{pi} cos(2t)+12cos(t)+19 dt\] \[2*(\frac{sin(2t)}{2}+12sin(t)+19t)\] \[sin(2t)+24sin(t)+38t;\ [0,pi]\]

OpenStudy (amistre64):

at 0 it all equals 0; so at pi it equals 38pi

OpenStudy (amistre64):

if i did it right; its been awhile :)